OpenStudy (anonymous):

1/3log(z)+5log(y)-log(x)

OpenStudy (solomonzelman):

What do you want to do here?

hartnn (hartnn):

express as a single logarithm

OpenStudy (anonymous):

+ multiplication - division

OpenStudy (solomonzelman):

at first, 1/3 will go inside as an exponent. and you will get z to the third, which is same as cube root of z. then apply your rules, \(\large\color{blue}{\log(a)+ \log(b)=\log(a\times b) }\) \(\large\color{blue}{\log(a)- \log(b)=\log(a\div b) }\)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np

hartnn (hartnn):

be careful with 5 log y too it will become log y^5

OpenStudy (solomonzelman):

yeah, I didn't see the 5, it will also go inside the log as an exponent of y

OpenStudy (solomonzelman):

I missed that...srry (Once was doing an integral of an absolute value, and kept getting it wrong and re-writing. My mistake as I figured out was that 1/2-1/4 is not 1/2, but 1/4 omg)

hartnn (hartnn):

LOL

OpenStudy (solomonzelman):

Paying attention to every step (as I failed to do just now) is very important!